OpenStudy (anonymous):

A normal distribution has a mean of 100 and a standard deviation of 12. Find the probability that a value selected at random is in the following interval. At least 76

OpenStudy (anonymous):

@triciaal can you help me out with this one?

OpenStudy (triciaal):

sorry not that fast with stats

OpenStudy (triciaal):

@campbell_st @Whitemonsterbunny17 @Squirrels do you mind helping?

OpenStudy (anonymous):

@Squirrels

OpenStudy (anonymous):

@Squirrels can you help me with this problem?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

@TylerD

OpenStudy (triciaal):

@jim_thompson5910

OpenStudy (kirbykirby):

\[P(X \ge 76)=P\left(\frac{X-100}{12}\ge\frac{76-100}{12}\right)=P(Z\ge -2) \]And this value you can find using a normal distribution table, a calculator/software , or by using the empirical rule (a.k.a. the 68–95–99.7 rule; knowing that __% of normally distributed date lies within 2 standard deviations of the mean).